I am confused on number 2 and I don’t know where to begin for 3b
e (PO4)2 3. a. Using the data given, calculate each of the thermodynamic quantibies for Balz dissolving in water at 25°C (use 298.15 K). Callculate and report each to 3 AH; in kl/mole 585.89 537.64 55.19 S in J/mole 165.14 9.62 AIC6213 AS060 AGe-L212 kl/mole reaction -kl/mole reaction b. From the ΔGe value, calculate the Ke for a saturated solution of Ball, Kap Type here to se DOLL
Answer #1

Use solubility rules to predict the solubility and insolubility in water.

insolubility in water:Fe3(PO4)2, AgBr, NiCO3 and CaSO4.

remaining given salts are soluble in water.

-------------------------------

formula we have to use is

\Delta G^{\circ}=-RTlnK_{eq}

here we can take keq =ksp

and rewriting the equation

k_{sp}=e^{\frac{-\Delta G^{\circ}}{RT}}

R =8.314 J.mol-1.K-1

T=298.15 K

\Delta G^{\circ}=-82.12 kJ/mol

k_{sp}=e^{\frac{-(-82.12)}{8.314\times 298.15}}

k_{sp}=e^{\frac{82.12}{2478.82}}

Ksp = e0.03312

Ksp = 1.03